OpenStudy (anonymous):

Solve for x: -2(x - 3) = 10 x = -2 x = 2 x = -8 x = 8

OpenStudy (anonymous):

option 1

OpenStudy (anonymous):

perhaps you should try these things yourself

OpenStudy (anonymous):

I am... you're the first person to actually give me the answer today. But thanks anyways.

OpenStudy (anonymous):

however, you know how to solve this?

OpenStudy (anonymous):

Um, I think, but I prefer help, I'm not good at math haha.

OpenStudy (anonymous):

divide both sides by -2, you will get, x-3=-5 then, x=-5+3=-2

OpenStudy (anonymous):

oh lol I would thought to do parenthesis first! lmao

OpenStudy (strawberry17):

Multiply -2 by both terms in the parentheses and then simplify: -2x + 6 = 10 -2x + 6 = 10 - 6 -6 ------------- -2x = 4 -2x 4 --- = --- -2 -2 x = -2

OpenStudy (anonymous):

alright thanks.

OpenStudy (strawberry17):

ur welcome.